A rectangle has the dimensions 10ft x 15ft. If it is dilated by a scale factor by1/5, what are the new dimensions?
Found 2 solutions by Aldorozos, ikleyn:
Answer by Aldorozos(172)

both dimensions have to be multiplied by 1/5
10*1/5 = 2
15*/15 = 3

Answer by ikleyn(52781)

.
both dimensions have to be multiplied by 1/5
10*1/5 = 2
15*1/5 = 3